# Inverse magneto-conductance design by automatic differentiation

### v1

@@ I Introduction
An alternative approach is to leverage automatic differentiation, which was recently developed by one of the authors[8, 9], to design Hamiltonians that reproduce target properties. Automatic differentiation allows us to solve the inverse problem without preparing training data, or relying on black box models such as neural networks. In addition, its computational cost remains manageable in practice, even when the number of parameters increases to, for example, a million. Previous studies have successfully designed systems with large quantum entanglement [9], a large anomalous Hall effect, and a photovoltaic effect [8]. Therefore, it can be used to identify microstructure that reproduces the target UCF. However, this approach has been applied primarily only to the diagonalization of the Hamiltonians of isolated systems. Application to dynamical properties in more general systems was left unexplored.
@@ II.1 Magneto-conductance calculation
In this study, quantum wires are modeled by a tight-binding model with a two-dimensional scattering region, and semi-infinite leads are connected at both ends. The effect of the magnetic field is taken into account as a Peierls phase with the Landau gauge. The Hamiltonian is given by
@@ III.2 Complicated Conductance Patterns
Next, we demonstrate that our framework can be applied to arbitrarily complex magneto-conductance patterns. As target magneto-conductance patterns, we have prepared two artificial patterns by tracing the shape of a Colosseum and Nirvana (a lying big Buddha), as indicated by the blue lines in Fig. 4(a) and (d). To allow for as flexible optimization as possible, the onsite potentials for each site, $h_{i}$, in the scattering region are treated as an independent parameter as $\bm{\theta}=(h_{1},h_{2},\dots,h_{900})\in\mathbb{R}^{900}$. The objective function $L(\bm{\theta})$ is the same as Eq. 4.
@@
Our results show that the framework can successfully reproduce the target patterns by optimizing the potential distributions. This highlights the versatility of our approach in handling various types of microscopic structures. We note that the solutions are not unique; during the numerical calculation, we have found several onsite potential distributions that reproduce the same magneto-conductance pattern by starting with different initial parameters. We also note that it is rather challenging to reproduce a pattern in which the conductivity increases when the magnetic field is increased because, in general, the conductivity decreases as the magnetic field due to the localization of electrons.
@@ III.3 Refining structure design
We have demonstrated that we can adjust the onsite potential distribution by adding a regularization term. This technique would reduce the need to design the potential distribution so precisely in an experimental setting. Further sophistication in the regularization term would allow for the design of more experimentally feasible potential distributions, which should be investigated in future works.
@@ IV Conclusion and outlook
Further investigation is required to improve our framework. Regarding defect identification, increasing the number of leads could facilitate the detection of more complex defects. In the current setup with only two leads, the defect position along the lead direction could not be identified due to translational symmetry; however, this issue can be mitigated by attaching leads in a different direction. Additional extensions would lead to detecting multiple defects and structural modifications. For designing magneto-conductance fluctuations, optimizing parameters in the Hamiltonian beyond the on-site potential, such as the hopping term, could enable more flexible structural designs. Moreover, the development of regularization terms, as discussed in Sec. III.3, will be crucial in future applications of our framework to experimental situations.

### v2

@@ I Introduction
An alternative approach is to leverage automatic differentiation, which was recently developed by one of the authors[8, 9], to design Hamiltonians that reproduce target properties. Automatic differentiation is a computational technique that computes the derivatives of a wide range of functions with analytical accuracy by systematically tracing each numerical operation. This method has been applied in neural network optimization, most notably in the backpropagation algorithm, by leveraging the gradient descent method. Automatic differentiation allows us to solve the inverse problem without preparing training data, or relying on black box models such as neural networks. In addition, its computational cost remains manageable in practice, even when the number of parameters increases to, for example, a million. Previous studies have successfully designed systems with large quantum entanglement [9], a large anomalous Hall effect, a photovoltaic effect [8], and a non-equilibrium Green function [10]. Therefore, it can be used to identify microstructure that reproduces the target UCF. However, this approach has been applied primarily only to the diagonalization of the Hamiltonians of isolated systems. Application to dynamical properties in more general systems was left unexplored.
@@ II.1 Magneto-conductance calculation
In this study, quantum wires are modeled by a tight-binding model with a two-dimensional scattering region, and semi-infinite leads are connected at both ends. Note that each site in the tight-binding model can be interpreted as a finite-difference discretization of a continuum Hamiltonian, where the model operates at a scale beyond the atomic level [11]. The effect of the magnetic field is taken into account as a Peierls phase with the Landau gauge. The Hamiltonian is given by
@@ III.2 Complicated Conductance Patterns
Next, we demonstrate that our framework can be applied to arbitrarily complex magneto-conductance patterns. As target magneto-conductance patterns, we have prepared two artificial patterns by tracing the shape of a Colosseum and Nirvana (a lying big Buddha), as indicated by the blue lines in Fig. 4(a) and (d). To allow for as flexible optimization as possible, the onsite potentials for each site, $h_{i}$, in the scattering region are treated as an independent parameter as $\bm{\theta}=(h_{1},h_{2},\dots,h_{900})\in\mathbb{R}^{900}$. The objective function $L(\bm{\theta})$ is the same as Eq. 4. The optimization of the onsite potentials corresponds to adjusting the gate voltages applied to the sample.
@@
Our results show that the framework can successfully reproduce the target patterns by optimizing the potential distributions. This highlights the versatility of our approach in handling various types of microscopic structures. We note that the solutions are not unique; during the numerical calculation, we have found several onsite potential distributions that reproduce the same magneto-conductance pattern by starting with different initial parameters. Different initial conditions yield different optimized parameters, but all of them reproduce the target conductance well. We also note that it is rather challenging to reproduce a pattern in which the conductivity increases when the magnetic field is increased because, in general, the conductivity decreases as the magnetic field due to the localization of electrons.
@@ III.3 Refining structure design
We have demonstrated that we can adjust the onsite potential distribution by adding a regularization term. When controlling the onsite potential using an electric field, a coarser structure, such as that on the right side of Fig. 5, is more feasible to realize than the finer structure on the left. Thus, this technique enables us to identify the on-site potential distributions that are easier to achieve among multiple solutions that reproduce magneto-conductance fluctuation. Further sophistication in the regularization term would allow for the design of more experimentally feasible ones, which should be investigated in future works.
@@ IV Conclusion and outlook
Further investigation is required to improve our framework. Regarding defect identification, increasing the number of leads could facilitate the detection of more complex defects. In the current setup with only two leads, the defect position along the lead direction could not be identified due to translational symmetry; however, this issue can be mitigated by attaching leads in a different direction. Additional extensions would lead to detecting multiple defects and structural modifications. For designing magneto-conductance fluctuations, optimizing parameters in the Hamiltonian beyond the on-site potential, such as the hopping term, could enable more flexible structural designs. Moreover, the development of regularization terms, as discussed in Sec. III.3, will be crucial in future applications of our framework to experimental situations. Additionally, it is important to address experimental uncertainties arising from impurities and other factors. One potential approach is to divide the process into two stages: first, constructing a tight-binding Hamiltonian through inverse analysis to reproduce the experimental results in the absence of a magnetic field, and then adjusting the on-site potential to replicate the desired magneto-conductance.